OpenStudy (anonymous):

Multiply (5-6x^6)^2 =

OpenStudy (anonymous):

This can be read as (5-6x^6)*(5-6x^6), right? So times out the terms; [5*5] + [5*-(6x^6)] + [5*-(6x^6)] + [(-6x^6)*(-6x^6)]

OpenStudy (anonymous):

We can then assess each term on it's own, if that's easier for you; 5*5 = 25 5*-(6x^6) = -30x^6 There are two of these factors, so now (-30x^6) + (-30x^6) = -60x^6 (-6x^6)*(-6x^6) = 36x^12 [remember your index laws! When multiplying two factors, add the powers: ie in this case, 6+6=12] Sum them up! 25 + -60x^6 + 36x^12 = 25 - 60x^6 + 36x^12